OpenStudy (anonymous):

Perfect Square Trinomial. Need Help Please. (x + 2y)2 @Hero Can you give me hand with this one as well? i keep getting it wrong?

hero (hero):

Apply the same process....the distributive rule a(b + c) = ab + ac You are given: (x + 2y)^2 = (x + 2y)(x + 2y) Treat the first (x + 2y) as a and the second (x + 2y) as (b+c) Then expand as follows: (x + 2y)(x + 2y) = x(x + 2y) + 2y(x + 2y) = x^2 + 2xy + 2xy + 4y^2 = x^2 + 4xy + 4y^2 This is a very systematic process

hero (hero):

It's a very systematic process, however, you still have to keep up with each operation.

OpenStudy (anonymous):

Okay thank you, this stuff is so confusing, it is the right answer, i just get lost trying to do the distributive rule, so thank you again, i've always been worst in math :p so dont be surprised if i may have a few more questions!

hero (hero):

Try to understand it for yourself first. Just make up your own problems and check the results yourself

OpenStudy (anonymous):

okay i will keep that in mind, and i will try my best to do so!